OpenStudy (anonymous):

Find the y-intercept of the line whose equation is 2x - 3y = -6.

OpenStudy (anonymous):

2x-3y=-6 +2x +2x -3y=2x-6 /-3 /-3 y=-2/3x+2

OpenStudy (anonymous):

-6 2 2/3 correct?

OpenStudy (anonymous):

2